Calculate the value of the g at a hight h = 4R from surface earth ( take g = 10m/s) at the earth


1
Expert's answer
2020-12-14T07:18:37-0500

The acceleration due to gravity on the surface of Earth can be found as follows:


"g=\dfrac{GM}{R^2}."


The acceleration due to gravity at a height "h" from surface Earth can be found as follows:


"g_{h}=\dfrac{GM}{(R+h)^2}."

Since, "h=4R" we get:


"g_{h}=\dfrac{GM}{(R+4R)^2}=\dfrac{GM}{(5R)^2}=\dfrac{1}{25}\dfrac{GM}{R^2}=\dfrac{g}{25},""g_{h}=\dfrac{10\ \dfrac{m}{s^2}}{25}=0.4\ \dfrac{m}{s^2}."

Answer:

"g_{h}=0.4\ \dfrac{m}{s^2}."